# A novel hybrid time-varying graph neural network for traffic flow forecasting

### v1

@@ Abstract
Real-time and accurate traffic flow prediction is the foundation for ensuring the efficient operation of intelligent transportation systems. In existing traffic flow prediction methods based on graph neural networks (GNNs), pre-defined graphs were usually used to describe the spatial correlations of different traffic nodes in urban road networks. However, the ability of pre-defined graphs used to describe spatial correlation was limited by prior knowledge and graph generation methods. Although time-varying graphs based on data-driven learning can partially overcome the drawbacks of pre-defined graphs, the learning ability of existing adaptive graphs was limited. For example, time-varying graphs cannot adequately capture the inherent spatial correlations in traffic flow data. In order to solve these problems, we have proposed a hybrid time-varying graph neural network (HTVGNN) for traffic flow prediction. Firstly, a novel time-aware multi-attention mechanism based on time-varying mask enhancement was reported to more accurately model the temporal dependencies among distinct traffic nodes in the traffic network. Secondly, we have proposed a novel graph learning strategy to concurrently learn both static and dynamic spatial associations between different traffic nodes in road networks. Meanwhile, in order to enhance the learning ability of time-varying graphs, a coupled graph learning mechanism was designed to couple the graphs learned at each time step. Finally, the effectiveness of the proposed method HTVGNN was demonstrated with four real data sets. Simulation results illustrate that, compared with the latest space-time graph neural network model, the proposed method HTVGNN has higher prediction accuracy. In addition, the ablation experiment verifies that the coupled graph learning mechanism can effectively improve the long-term prediction performance of HTVGNN. The source code is located at: https://github.com/DaibenaoZSTU/HTVGNN.
@@
Nowadays, with the acceleration of the urbanization process, the car ownership has increased sharply. Subsequently, how to improve the traffic efficiency of the road network to alleviate traffic congestion has always been a difficult problem [1]. Since it can provide the traffic state information for a certain period of time in the future based on real-time monitored traffic data, traffic flow prediction has been widely applied to improve the level of intelligent traffic management and control in road networks. In the early studies, researchers mainly used mathematical statistics based methods to construct traffic flow prediction models. For example, historical average (HA) [2] [3], autoregressive moving average model (ARIMA) [4] [5], vector autoregressive model (VAR) [6], etc. These statistics based methods assume that the dependency of time-series data is linear. However, since traffic flow often has complex and changeable characteristics, the dependency of the traffic flow time-series data is nonlinear. As a result, the performance of statistics based methods is not well in actual traffic flow prediction tasks. After that, traditional machine learning-based methods (e.g. support vector regression (SVR) [7] and k-nearest neighbor (KNN) [8]) were reported to overcome the drawbacks of statistics based methods. Nevertheless, the performance of traditional machine learning-based methods mainly relies on the handcrafted features, which makes it difficult for them to fully utilize the increasingly abundant large-scale traffic flow data. Then, with the rapid development of deep learning, long short-term memory (LSTM) and their variants gated recurrent units (GRU) [9] were used to model the temporal dependencies of traffic flow time-series data, while convolutional neural networks (CNN) [10] and graph convolutional networks (GCN) [11] [12] [13] were employed to capture the spatial dependencies of different traffic nodes.
The first challenge lies in capturing the spatial correlation between roads more effectively, without relying on prior knowledge. Previous studies have predominantly relied on certain prior knowledge (such as distance, DTW, Pearson coefficient) to construct a spatial structure map that signifies the spatial relationship of road networks. Examples include Spatiotemporal Graph Convolution (STGCN) [14] and Spatiotemporal Fusion Graph Neural Network (STFGNN) [15], which calculate node correlations based on road distances and historical traffic data to establish a predefined spatial structure map between roads. However, these predefined maps often fail to accurately represent the traffic flow data and our prediction model. The expressiveness of such predefined maps depends heavily on the rationality of prior knowledge and graph generation methods employed, resulting in potential inclusion of unreasonable or incorrect spatial correlations that limit their effectiveness. To address this issue, Graph WaveNet (Graph WaveNet) has achieved some success by utilizing predetermined graphs for both forward and backward diffusion while incorporating adaptive graphs. Nevertheless, GWN still heavily relies on predetermined graphs for capturing spatial correlation. Adaptive Graph Convolutional Recurrent Network (AGCRN) [16] and Structure-aware Convolutional Neural Networks (SACNNs) [17] have further improved adaptive spatial modeling by achieving similar effects as predefined graph structures without any reliance on predetermined graphs at all.However, current adaptive graph learning typically involves learning a global adaptive graph for all time steps, thereby limiting its learning capability.
Furthermore, regardless of whether it is an adaptive or predefined graph, the spatial correlation described by it remains constant. However, the traffic network exhibits complexity and variability, with its topology undergoing intricate changes over time. Effectively modeling the dynamic spatial correlation between roads is therefore crucial in spatial correlation modeling.
@@
In this paper, we propose a novel time-aware multi-head attention mechanism based on time-varying mask enhancement to address the aforementioned issues. The aim is to enhance the computational accuracy and time-awareness capability of multi-head attention.we focus on capturing the spatio-temporal characteristics of complex changes by introducing an efficient coupled time-varying graph convolutional recurrent network. This network utilizes time-varying graphs at each time step to learn the spatial dependencies within the road network. to improve its effectiveness in learning long-term dependencies, we integrate the learned time-varying graphs from different time steps. This approach enhances prediction performance by considering short-term traffic patterns between adjacent time steps. additionally, for effectively modeling dynamic spatial correlation within the road network, we propose a dynamic time-varying graph convolution that combines topology matrix and similarity matrix as a mask matrix for more effective modeling of dynamic spatial correlations.
The following are the four key contributions of our work in traffic flow prediction:
A novel time-aware multi-head attention mechanism based on time-varying mask enhancement was proposed. Different from the traditional multi-head attention mechanism, the proposed time-aware multi-head attention mechanism can allocate the optimal attention calculation scheme according to the time characteristics of the input data, so as to more accurately model the time dependence relationship between different traffic nodes in the road network. The time-varying mask was composed of two parts: the static mask was embedded to correct the multi-head attention, and the dynamic mask was embedded to enhance the time-sensing ability of the time-head attention.
A novel graph learning strategy was proposed, which can learn static spatial association and dynamic spatial association in parallel. In order to learn the static spatial correlation better, the defined global embedding vector and local embedding vector were used to learn the spatial correlation of the road network at each time step, and a coupled graph learning mechanism was designed to couple the graphs learned at each time step. In order to model the dynamic spatial correlation in the road network more accurately, a mask based on topological matrix and semantic matrix was defined for the first time. Meanwhile, the mask was used to optimize the dynamic graph generated by the graph learning strategy.
@@
Simulation test results based on four real data sets verify the effectiveness of the proposed method HTVGNN. Compared with the latest space-time graph neural network model, the proposed HTVGNN algorithm has higher prediction accuracy. In addition, the ablation experiment verifies that the coupled graph learning mechanism in the proposed graph learning strategy can effectively improve the long-term prediction performance of the model.
@@ II-A Traditional Traffic Forecasting Methods
Traffic flow prediction plays a pivotal role in urban intelligent transportation, enabling the anticipation of future traffic conditions based on real-time data from the traffic network. This information serves as a foundation for scheduling and decision-making by the traffic management department [18]. In earlier studies, researchers predominantly relied on mathematical statistical methods and early machine learning techniques, such as historical average (HA) [12] [11], autoregressive integrated moving average (ARIMA) [10] [11], vector autoregressive model (VAR) [19], etc. However, these models often underperform in actual traffic prediction tasks due to the complex and dynamic nature of traffic flow. With significant advancements in deep learning across various domains like speech recognition and image processing, an increasing number of researchers have started employing deep learning approaches for spatio-temporal traffic flow prediction tasks. Convolutional Neural Networks (CNN) [10], Long Short-Term Memory (LSTM) networks [9] along with its variant Gate Recurrent Unit (GRU) [9] are widely used and representative deep learning methods that have significantly enhanced the accuracy of traffic flow prediction. Unlike previous methods, deep learning-based approaches can effectively capture spatio-temporal correlations. For instance, Fully Connected Long Short-Term Memory Network (FC-LSTM) [20] combines CNN and LSTM to extract implicit spatio-temporal characteristics present in traffic flow data. Nevertheless, since CNN was originally designed to process Euclidean data with regular grid structures like images and videos rather than road networks’ spatial features accurately; FC-LSTM’s predictive performance is suboptimal.
@@ II-C Transformer
The Transformer model was initially proposed in the field of machine translation [23] and has since been widely adopted in deep learning domains, including NLP [24] and computer vision [25]. It is a codec structure implemented with multi-head self-attention (MSA), which exhibits strong modeling capabilities for contextual information. Consequently, MSA has found extensive applications in machine translation tasks, delivering superior performance. In the context of traffic flow prediction tasks, the Transformer model incorporates a time-based multi-head attention mechanism that maps data to high-dimensional space using multiple groups of attention mechanisms. This enables simultaneous learning of optimal attention between different features. Researchers have successfully introduced MSA into traffic flow prediction, yielding significant improvements across various aspects. For instance, Yan et al. [26] proposed the Traffic Transformer model by first employing a designed coding and embedding block to address data discrepancies across different research fields. Additionally, they enhanced the original encoder-decoder structure of Transformer by introducing two components: global encoder and global-local decoder blocks. These blocks are stacked together to form a deep prediction network with hierarchical characteristics. Wang et al.[27], on the other hand, presented a novel graph neural network layer incorporating position-wise attention mechanism that dynamically aggregates historical traffic flow information from adjacent roads. To better capture temporal features at both local and global levels, they combine RNN with a Transformer layer.
III Methodology
III-A Problem defination
@@
Definition 1. The traffic network can be represented as a weighted directed graph , where represents all nodes (sensors) in the road network ) (), is the set of all edges, which represents the connectivity between nodes. represents the weighted adjacency matrix, and its value represents the connectivity between nodes. degree of correlation.
@@
Definition 2. Based on the transportation network, we define the problem of traffic flow prediction with input represents the traffic flow characteristics observed by all nodes at a certain time (such as traffic flow, speed, etc.)
@@
Definition 3. Given the road topology and traffic characteristics . Our goal is to predict the future traffic conditions at times by analyzing the data at times in the past.
Definition 4. If there are nodes in the traffic network, and the sensor in the network sampling d times per day. The week consists of a total of 7 days. Then the static mask embedding and dynamic mask embedding can be represented by three independent trainable sub-mask embeddings and , . represents the dimension of sub-mask embedding.
IV Hybrid time-varying graph neural network
@@
The present study proposes a time trend-aware multi-head self-attention transformer and a graph convolution GRU-based traffic flow prediction model. The former effectively captures local temporal features, while the latter simultaneously extracts spatio-temporal features. Firstly, we introduce the fundamental structure of HTVGNN and illustrate it in Fig. 1. To better represent intricate correlations among nodes, we propose static and dynamic time-varing graphs that learn changing spatial correlations through trainable node embeddings. Subsequently, graph convolution GRU is employed to extract spatio-temporal features for accurate traffic flow prediction. Detailed model structures and mechanisms are presented in subsequent subsections.
@@ IV-A An enhanced temporal perception multi-head self-attention
Self-attention mechanism is a broader attention mechanism where query, key and value are different representations of the same sequence. In order to learn the complex and time-varying temporal dependencies of input data, the temporal feature is usually mapped to a high-dimensional space with a widely used multi-head self-attention (MHSA) [1]. For the self-attention head at time step t, and are generated by the linear projection as defined in Eq. (1), respectively.
@@
Finally, the final multi-head self-attention corresponding to the input can be obtained by weighting the attention of the self-attention heads together. Specifically, it can be calculated by the following Eq. (3),
where is the total amount of attention heads. is the final output projection matrix. However, in traditional multi-head attention mechanism, the Key vectors () will cause unnecessary interference [Weiguo Zhu]. In addition, since the learned matrix , , and are static, the score calculated by the traditional multi-head attention equation defined in Eq. (2) can not fully perceive the temporal trend of self-attention. Therefore, we have proposed an enhanced temporal perception multi-head self-attention. For notational simplicity, our proposed enhanced temporal perception multi-head self-attention (ETPMSA) as shown in Fig.2, ETPMSA can be calculted by the following as Eq.(4):
where is a learnable parameter vector, denotes the self-attention head which can be calculted by the following Eq. (5).
where and are the Query, Key and Value of the traditional time multi-head attention as defined in Eq. (2), is adaptive mask matrix which is defined according to the size of the input data. It should be noted that, is the mask embedding which composed by three submask embedding , , . Herein, submask embedding is defined to enhance the accuracy of multi-head self-attention. submask embeding , are daily mask embeding and weekly mask embedding whose time step is constitent with the history data .
IV-B Time-varying graph convolution recurrent module
The architecture of our proposed TVGCRM is illustrated in Fig. 3. The details of the TVGCRM is summarized as follows.
IV-B1 Graph learning
There usually exist static and dynamic spatial correlations between different traffic nodes. In order to capture the static and dynamic spatial correlations, we have proposed a graph learning method to generate static graph and dynamic graph in a data-driven manner without providing any prior knowledge.

### v3

@@ Abstract
Real-time and precise traffic flow prediction is vital for the efficiency of intelligent transportation systems. Traditional methods often employ graph neural networks (GNNs) with pre-defined graphs to describe spatial correlations among traffic nodes in urban road networks. However, these pre-defined graphs are limited by existing knowledge and graph generation methodologies, offering an incomplete picture of spatial correlations. While time-varying graphs based on data-driven learning have attempted to address these limitations, they still struggle with adequately capturing the inherent spatial correlations in traffic data. Moreover, most current methods for capturing dynamic temporal correlations rely on a unified calculation scheme using a temporal multi-head self-attention mechanism, which often leads to inaccuracies. In order to overcome these challenges, we have proposed a novel hybrid time-varying graph neural network (HTVGNN) for traffic flow prediction. Firstly, a novel time-aware multi-attention mechanism based on time-varying mask enhancement was reported to more accurately model the dynamic temporal dependencies among distinct traffic nodes in the traffic network. Secondly, we have proposed a novel graph learning strategy to concurrently learn both static and dynamic spatial associations between different traffic nodes in road networks. Meanwhile, in order to enhance the learning ability of time-varying graphs, a coupled graph learning mechanism was designed to couple the graphs learned at each time step. Finally, the effectiveness of the proposed method HTVGNN was demonstrated with four real data sets. Simulation results revealed that HTVGNN achieves superior prediction accuracy compared to the state of the art space-time graph neural network models. Additionally, the ablation experiment verifies that the coupled graph learning mechanism can effectively improve the long-term prediction performance of HTVGNN.
@@
The rapid acceleration of urbanization has led to a sharp increase in car ownership. This rise in vehicles has intensified the challenge of enhancing road network efficiency to alleviate traffic congestion [1]. Traffic flow prediction, offering future traffic state information based on real-time data, is pivotal for advancing intelligent traffic management and control.
@@
Early research primarily utilized statistical methods for constructing traffic flow prediction models. For example, historical average (HA) [2] [3], autoregressive moving average model (ARIMA) [4] [5], vector autoregressive model (VAR) [6], etc. These statistics based methods assume linear dependencies in time-series data. However, the complex and variable nature of traffic flow leads to nonlinear dependencies, resulting in these methods’ subpar performance in practical forecasting tasks. After that, traditional machine learning-based methods (e.g. support vector regression (SVR) [7] and k-nearest neighbor (KNN) [8]) were reported to overcome the drawbacks of statistics based methods. Nevertheless, such methods’ reliance on handcrafted features limits their effectiveness in leveraging the growing abundance of large-scale traffic flow data. After that, long short-term memory (LSTM) and their variants gated recurrent units (GRU) [9] were used to model the temporal dependencies of traffic flow time-series data, and convolutional neural networks (CNN) [10] and graph convolutional networks (GCN) [11] [12] [13] were employed to capture the spatial dependencies of different traffic nodes.
In recent years, since the core of traffic flow prediction is to accurately model the complex and nonlinear spatial-temporal dependencies of dynamic traffic flows of different traffic nodes, graph convolutional network (GCN) and Transformer [14] are two widely used core components of many representative traffic flow prediction models [15, 16, 17, 18]. While such spatial-temporal modeling approaches have yielded promising results in traffic flow prediction, accurately capturing the complex and dynamic spatial-temporal correlations within road networks remains challenging for the following reasons.
Initially developed for natural language processing, the Transformer model[14] excels in machine translation due to its strong ability to model long-term contextual features. Its Encoder-Decoder architecture notably improves model generalization. Central to this architecture is the multi-head self-attention mechanism. However, this traditional mechanism struggles to effectively model dynamic temporal correlations in traffic flow, often leading to errors in weight calculations [17]. Additionally, integrating periodic features into this mechanism could potentially enhance its capacity to perceive time-related variations in traffic data.
It is challenging to capture the spatial correlations between road network nodes without prior knowledge. Although some works try to create spatial structure maps based on existing knowledge [19, 20], these maps often inaccurately represent spatial correlations due to the limitations of their underlying assumptions and methods. Some other works [21, 22] have made strides in adaptive spatial modeling without fixed graphs, but they still predominantly use a global adaptive graph across all time steps, restricting their learning potential.
Additionally, a major challenge lies in the static nature of spatial correlations in both adaptive and predefined graphs, which doesn’t align with the dynamic and ever-changing topology of traffic networks. Accurately modeling these dynamic spatial correlations is crucial. Existing models like Liu et al.’s STGHTN [16] and Zhu et al.’s CorrSTN [18] use multi-head attention mechanisms for dynamic modeling-the former for adjusting features dynamically and the latter for modifying graphs before cluster convolution. However, these approaches typically entail high computational complexity due to their reliance on spatial multi-head attention.
@@
To tackle the identified challenges, this paper introduces a novel time-aware multi-head attention mechanism enhanced by time-varying masks, aiming to boost the computational accuracy and temporal sensitivity of attention models. We present an efficient coupled time-varying graph convolutional recurrent network, designed to capture the spatio-temporal dynamics of traffic flows. This network leverages time-varying graphs for each time step, enhancing its ability to learn spatial dependencies and long-term trends in road network traffic. By integrating time-varying graphs across different time steps, our model effectively incorporates both short-term and long-term traffic patterns. Furthermore, we propose a dynamic graph convolution approach, which utilizes a combined topology and similarity matrix as a mask, enabling more precise modeling of the dynamic spatial correlations in traffic networks.
The following are three key contributions of our work in traffic flow prediction:
A novel temporal perception multi-head self-attention mechanism enhanced by a time-varying mask, diverging from traditional approaches is proposed. This mechanism dynamically adjusts attention calculations based on the temporal characteristics of input data, enabling more accurate capture of temporal dependencies among traffic nodes. The time-varying mask consists of three components: a static mask embedding for correcting the multi-head self-attention mechanism, and a dynamic mask embedding to augment the time-awareness of the mechanism.
A new graph learning strategy is introduced to simultaneously address static and dynamic spatial correlations in traffic networks. To enhance learning of static spatial correlations, we employ both global and local time-varying embeddings, facilitating the understanding of spatial relationships at various time steps. Additionally, a coupling graph learning mechanism is implemented to integrate graphs learned at different time steps. For more precise capture of dynamic spatial correlations, we define a mask matrix, derived from topology and semantic matrices, to refine the dynamic graphs produced by our strategy.
@@
Simulation on four real-world datasets is made to evaluate our proposed HTVGNN method. The results demonstrate the effectiveness of HTVGNN, showing higher prediction accuracy than the latest space-time graph neural network models. What’s more, an ablation study confirms that the integrated coupled graph learning mechanism significantly enhances the model’s long-term prediction capabilities.
@@ II-A Traditional Traffic Forecasting Methods
Traffic flow prediction plays a pivotal role in urban intelligent transportation, enabling the anticipation of future traffic conditions based on real-time data from the traffic network. This information serves as a foundation for scheduling and decision-making by the traffic management department [23]. In earlier studies, researchers predominantly relied on mathematical statistical methods and early machine learning techniques, such as historical average (HA) [12] [11], autoregressive integrated moving average (ARIMA) [10] [11], vector autoregressive model (VAR) [19], etc. However, these models often underperform in actual traffic prediction tasks due to the complex and dynamic nature of traffic flow. With significant advancements in deep learning across various domains like speech recognition and image processing, an increasing number of researchers have started employing deep learning approaches for spatio-temporal traffic flow prediction tasks. Convolutional Neural Networks (CNN) [10], Long Short-Term Memory (LSTM) networks [9] along with their variant Gate Recurrent Unit (GRU) [9] are widely used and representative deep learning methods that have significantly enhanced the accuracy of traffic flow prediction. Unlike previous methods, deep learning-based approaches can effectively capture spatio-temporal correlations. For instance, Fully Connected Long Short-Term Memory Network (FC-LSTM) [15] combines CNN and LSTM to extract implicit spatio-temporal characteristics present in traffic flow data. Nevertheless, CNNs, originally designed for Euclidean data with regular grid structures such as images and videos, do not accurately process the spatial features of road networks, leading to suboptimal predictive performance in FC-LSTM.
@@ II-C Transformer
The Transformer model was initially proposed in the field of machine translation [14] and has since been widely adopted in deep learning domains, including NLP [26] and computer vision [27].It features a codec structure implemented with multi-head self-attention (MSA), known for its robust modeling of contextual information. Consequently, MSA has found extensive applications in machine translation tasks, delivering superior performance. For traffic flow prediction, the Transformer model employs a time-based multi-head attention mechanism, mapping data to high-dimensional space through multiple attention groups. This enables simultaneous learning of optimal attention between different features. Researchers have successfully introduced MSA into traffic flow prediction, yielding significant improvements across various aspects. For instance, Yan et al. [28] proposed the Traffic Transformer model by first employing a designed coding and embedding block to address data discrepancies across different research fields. Additionally, they enhanced the original encoder-decoder structure of Transformer by introducing two components: global encoder and global-local decoder blocks. These blocks are stacked together to form a deep prediction network with hierarchical characteristics. Wang et al.[29], on the other hand, presented a novel graph neural network layer incorporating position-wise attention mechanism that dynamically aggregates historical traffic flow information from adjacent roads. To better capture temporal features at both local and global levels, they combine RNN with a Transformer layer.
III Problem Definition
@@
The traffic network can be represented as a weighted directed graph , where represents all nodes in the road network (). is the set of all edges, which represents the connectivity between different nodes. represents the weighted adjacency matrix, and its value represents the degree of the correlation between different nodes.
@@
Based on the traffic network, the input of the traffic flow prediction is defined with , which represents the traffic flow characteristics (such as traffic flow, speed, etc.) observed by all nodes at times in the past.
@@
Given the road topology and traffic flow characteristics at times in the past, our goal is to predict the traffic conditions at times in the future by analyzing the traffic flow characteristics .
In the traffic network with nodes, where sensors sample data times a day ( = 288 if the sampling interval is set to minutes) and considering a week of days. The dimension of the embedding is . Define three independently trainable embeddings , and . and are used to find and generate the daily sub-mask embedding and the weekly sub-mask embedding corresponding to the current traffic flow . Meanwhile, is designed to optimize the generated daily sub-mask embedding and weekly sub-mask embedding .
IV Methodology
@@
The present study proposes a time trend-aware multi-head self-attention transformer and a graph convolution GRU-based traffic flow prediction model. The former effectively captures local temporal features, while the latter simultaneously extracts spatio-temporal features. Firstly, we introduce the fundamental structure of HTVGNN and illustrate it in Fig. 1. To better represent intricate correlations among nodes, we propose static and dynamic time-varying graphs that learn changing spatial correlations through trainable node embeddings. Subsequently, graph convolution GRU is employed to extract spatio-temporal features for accurate traffic flow prediction. Detailed model structures and mechanisms are presented in subsequent subsections.
@@ IV-A An enhanced temporal perception multi-head self-attention
Self-attention mechanism is a broader attention mechanism where query, key and value are different representations of the same sequence. In order to learn the complex and time-varying temporal dependencies of input data, the temporal feature is usually mapped to a high-dimensional space with a widely used multi-head self-attention (MHSA). For the head self-attention, and are generated by the linear projection as defined in Eq.(1), respectively.
@@
Finally, the final multi-head self-attention corresponding to the input can be obtained by weighting the attention of the self-attention heads together. Specifically, it can be calculated by the following equation (3),
where is the total amount of attention heads. is the final output projection matrix. However, in traditional multi-head attention mechanism, the Key vectors () will cause unnecessary interference [17]. In addition, since the learned matrix , , and remained unchanged after training, the score calculated by the traditional multi-head attention equation defined in Eq. (2) can not fully perceive the temporal trend of self-attention. Therefore, a time-varying mask embedding generator is designed. The key process of the generator finds the daily submask embedding and the weekly submask embedding of the time corresponding to the current traffic flow . Then, based on an independent trainable embedding , we have proposed an enhanced temporal perception multi-head self-attention. For notational simplicity, our proposed enhanced temporal perception multi-head self-attention (ETPMSA), as shown in Fig.2, can be calculated as follows in Eq.(4):
where is a learnable parameter vector, denotes the self-attention head which can be calculted by the following Eq.(5) to Eq.(9).
where and are the Query, Key and Value of the traditional time multi-head attention as defined in equation (2), is mask matrix which is defined according to the size of the input data at the time step . It should be noted that, is the mask embedding which composed of three submask embedding and , . Submask embedding and are the daily submask embedding and weekly submask embedding at . The submask embedding and are employed to enhance the temporal perception by updating the key of multi-heads self-attention mechanism.
IV-B Position Encodings
Among them, represents the position and represents the feature dimension, denotes the dimension of the model.
IV-C Coupling time-varying graph convolution recurrent module
In order to further extract the temporal and spatial characteristics, a novel coupling time-varying graph convolution recurrent module (CTVGCRM) is proposed. The architecture of our proposed CTVGCRM is illustrated in Fig. 3. The details of the CTVGCRM are summarized as follows.
IV-C1 Graph learning
As pointed out in [30] [31], there usually exist static and dynamic spatial correlations between different traffic nodes. In order to capture the static and dynamic spatial correlations, we have proposed a graph learning method to generate static graph and dynamic graph in a data-driven manner without providing any prior knowledge.